A roofer and an assistant can repair a roof together in 8 hours. Working alone, the assistant can complete the repair in 17 hours. If both the roofer and the assistant work together for 7 hours and then the assistant is left alone to finish the job, how much longer should the assistant need to finish the repairs?
Answer by lwsshak3(11628)

A roofer and an assistant can repair a roof together in 8 hours. Working alone, the assistant can complete the repair in 17 hours. If both the roofer and the assistant work together for 7 hours and then the assistant is left alone to finish the job, how much longer should the assistant need to finish the repairs?
**
If roofer and assistant worked together for 7 hours, the completed 7/8 of the job, leaving 1/8 of the job for the assistant to finish.
let x=hours assistant must work to finish the repairs
time=work/rate
x=(1/8)/(1/17)=17/8 hrs